OpenStudy (anonymous):

A particle is moving along the curve y=4 times square root of (4x+5).As the particle passes through the point (5, 20), its x-coordinate increases at a rate of 3 units per second. Find the rate of change of the distance from the particle to the origin at this instant.

OpenStudy (anonymous):

Where did you get 0 from?

OpenStudy (anonymous):

Why was the coordinate point given? Extra information?

OpenStudy (anonymous):

am I plugging in 5 for x?

OpenStudy (anonymous):

The 0's come from the origin, \((0,0)\). Review the distance formula if you're still not sure. The coordinate point gives you everything you need. However, if you were to use the y-coordinate you would have needed some information about the rate of change of that coordinate. You weren't, so we improvise; instead of using an exact number, we use the given equation. And yes, \(x=5\) and \(\dfrac{dx}{dt}=3\).

OpenStudy (anonymous):

I am getting 27/250. It is coming out wrong

OpenStudy (anonymous):

Oops, tiny mistake. \[y=4\sqrt{4x+5}\] Let \((x,y)\) be a point that travels along the curve. The point will be equivalent to \(\left(x,4\sqrt{4x+5}\right)\). The distance between any point on the curve and the origin will be \[D=\sqrt{(x-0)^2+(y-0)^2}=\sqrt{x^2+16(4x+5)}=\sqrt{x^2+64x+80}\] Now do some implicit differentiation: \[\frac{dD}{dt}=\frac{1}{2}\left(x^2+64x+80\right)^{-1/2}\left(2x+64\right)\frac{dx}{dt}\] Plug in your given info.

OpenStudy (anonymous):

Now it worked